# First investigation of void statistics in numerical relativity simulations

### v1

@@ Abstract
We apply and extend standard tools for void statistics to cosmological simulations that solve Einstein’s equations with numerical relativity (NR). We obtain a simulated void catalogue without Newtonian approximations, using a new watershed void finder which operates on fluid-based NR simulations produced with the Einstein Toolkit. We compare and contrast measures of void size and void fraction, and compare radial stacked density profiles to empirically-derived Hamaus–Sutter–Wandelt (HSW) density profiles and profiles based on distance to void boundaries. We recover statistics roughly consistent with Newtonian $N$-body simulations where such a comparison is meaningful. We study variation of dynamical spatial curvature and local expansion explicitly demonstrating the spatial fluctuations of these quantities in void regions. We find that voids in our simulations expand $\sim$10–30faster than the global average and the spatial curvature density parameter in the centre of voids reaches $\sim$60–80%. Within the limits of resolution of the simulations, the results are consistent with the parameters of the Timescape model of cosmological backreaction.
@@
cosmology: theory – gravitation – large-scale structure of Universe – methods: data analysis
@@ 1 Introduction
Observations present us with the immediate dichotomy that voids – the absence of matter – by their very nature can only be understood by directly observing the complex matter distribution within which they grow and interact. Observations need to be combined with modelling and numerical simulations to glean any understanding. The goal of this paper is to extend that modelling to a full general relativistic analysis of void statistics for the first time.
@@
Simulated void catalogues are typically produced by Newtonian $N$-body simulations, in which expansion is prescribed by scaling the simulation volume according to the Friedmann equation from an FLRW model. However, in GR matter and space-time curvature are non-linearly related. Consequently, regional expansion in the universe is determined by its matter content, allowing over-dense and under-dense regions to have varying degrees of spatial curvature. This may occur even if the average spatial curvature of the Universe is zero (as in $\Lambda$CDM). This is particularly relevant for cosmic voids, which have a negative spatial curvature and expand at a greater rate than the over-dense sheets, filaments and knots of the cosmic web (Peebles, 1993, 2001; Macpherson et al., 2018). Due to the detachment of matter and space-time in Newtonian gravity, it has not been possible to directly capture these effects with traditional $N$-body simulations.
Traditional $N$-body codes have been been significantly improved, most notably by gevolution (Adamek et al., 2016) which uses the weak-field limit of GR. In simplified setups, gevolution has been shown to agree with our full numerical relativity approach (Adamek et al., 2020, 2024). However, given inherent technical ambiguities with the choice of a global background in any weak field GR approach, non-linear GR might be desirable in some scenarios. Another alternative to evolving all degrees of freedom contained in the Einstein equations is to self–consistently remove degrees of freedom with likely small contributions in cosmology: (i) gravitational waves in the gramses (Barrera-Hinojosa & Li, 2020a, b) code; (ii) vorticity and the magnetic Weyl curvature in quiet universes (Heinesen & Macpherson, 2022); and (iii) all of the above in silent universes (Bolejko, 2018a). Silent universe models have been used to seek solution to the Hubble tension (Bolejko, 2018b). However, the relationship between the Bolejko (2018b) solution and realistic initial conditions from the Cosmic Microwave Background (CMB) epoch are not clear. We aim to circumvent such limitations.
@@ 1.2 Backreaction of inhomogeneities
Different types of backreaction have been classified in the well-known Buchert et al. (2020) (BMR) formalism, which provides a set of equations for the evolution of scalar averages of small scale structures. The first Buchert equation may be written as the energy density sum-rule:
$\bar{\Omega}_{\mathrm{M}}+\bar{\Omega}_{\mathrm{R}}+\bar{\Omega}_{\mathrm{K}}+\bar{\Omega}_{\mathcal{Q}}=1\,.$ (1)
@@
Here $\bar{\Omega}_{\mathrm{M}}$ and $\bar{\Omega}_{\mathrm{R}}$ are analogous to matter and radiation energy densities in the Friedmann equation, but scale with respect to powers of an average volume scale $\bar{a}$, (cf. (11) below), not a background metric scale factor; $\bar{\Omega}_{\mathrm{K}}$ is the average spatial curvature or kinetic curvature density; and $\bar{\Omega}_{\mathcal{Q}}$ is the kinematical backreaction density [^1] A cosmological constant backreaction density, $\bar{\Omega}_{\Lambda}$, and additional backreaction terms are also found in general (BMR), but are not relevant for interpreting our simulations. The radiation density $\bar{\Omega}_{\mathrm{R}}$ is not relevant except insofar as the ratio $\bar{\Omega}_{\mathrm{R}}/\bar{\Omega}_{\mathrm{M}}\simeq 0.3$ at last scattering, with significance for calibrations relative to CMB data. Our notation differs from that of BMR, who use $\Omega_{\mathcal{R}}$ in place of $\bar{\Omega}_{\mathrm{K}}$, while not explicitly writing a radiation density. Positive values $\bar{\Omega}_{\mathrm{K}}>0$ correspond to the negative spatial curvature, as is the case for void domination. .
[…]
By contrast, a positive kinetic curvature term $\bar{\Omega}_{\mathrm{K}}$ makes a very significant contribution to the late epoch energy budget (1) in several backreaction models, including those with an appreciable cosmological constant (Bolejko, 2018b). In the Timescape model $\bar{\Omega}_{\mathrm{K}}\propto f_{\rm v}^{1/3}/(\bar{a}^{2}\bar{H}^{2})$ has a direct physical interpretation as the kinetic energy of expansion of voids, [^2] As well as ensuring that $\bar{\Omega}_{\mathrm{K}}\to 0$ at early times when $f_{\rm v}\to 0$, the void volume fraction, $f_{\rm v}$, is the only free function, as $\bar{\Omega}_{\mathrm{M}}$ is also related to it (TS). completely replacing dark energy.
@@
Just as forces are produced by gradients of Newtonian potentials, in GR one deals with single derivatives of the connection, or two derivatives of the 10 metric coefficients. Likewise the gradient of the kinetic curvature, $\bar{\Omega}_{\mathrm{K}}$, sources dynamical spatial curvature variation. Inhomogeneities are the source of local and quasilocal (i.e., regional) anisotropies in both GR and its weak field Newtonian limit. However, generically in GR such anisotropies are not limited to a global FLRW background plus the spherical multipole expansion obtained by local Lorentz boosts in $v/c$ power series. The differences have been classified via multipole moments as nonkinematic differential expansion (Bolejko et al., 2016; McKay & Wiltshire, 2016; Dam, 2016).
@@
The methods of Section 1.1 have enabled the first numerical investigations into the impact of backreaction in late-time cosmology (Bentivegna & Bruni, 2016; Macpherson et al., 2018, 2019)and the study of cosmological observables in a non-linear, general-relativistic framework (Giblin et al., 2017; Macpherson, 2023). The characteristic features of quasilocal anisotropies expected in GR, as constrained by the initial matter power spectrum, are an important fundamental question. This paper quantifies these issues for the first time.
@@ 2 Numerical Relativity Simulations
We evolve the simulations from the initial redshift through to $z\approx 0$ using the 3+1 Baumgarte–Shapiro–Shibata–Nakamura–Oohara–Kojima (BSSNOK or BSSN, Baumgarte & Shapiro, 1998; Shibata & Nakamura, 1995; Nakamura, Oohara & Kojima, 1987) formalism of NR. The final slice of the simulation is defined based on the predicted change in volume from the initial background EdS model. The space-time is evolved using the McLachlan (Brown et al., 2009) ML_BSSN thorn in the ET and the hydrodynamics is evolved using GRHydro (Baiotti et al., 2005). We approximate matter as dust, namely, as a pressureless [^5] GRHydro does not allow for an exactly zero pressure, so we choose $P\ll\rho$ using a Polytropic equation of state (EOS), which is sufficient to approximate dust (Macpherson et al., 2017). Such an EOS has also been shown to be beneficial in analytic studies in inhomogeneous cosmology (Bolejko & Lasky, 2008). perfect fluid (under a continuous fluid approximation). See Macpherson et al. (2017) and Macpherson et al. (2019) for further technical specifics of the simulation set up.
@@ 3 Watershed Void Finder
This method originates with the Watershed Void Finder (Platen, van de Weygaert & Jones, 2007) and zobov (Neyrinck, 2008). One popular watershed void finder based on zobov is vide (Sutter et al., 2014a), which has been used in the study of both observational data (e.g., Sutter et al., 2012, 2014b) and simulations (e.g., Contarini et al., 2022).
Most void-finding software is designed to operate on sets of particle positions, either from observational data (positions of galaxies) or the output of $N$-body simulations (Platen et al., 2007; Neyrinck, 2008; Sutter et al., 2014a). For watershed void-finding, it is necessary to translate these into a continuous density field upon which the watershed transform can be performed. It is typical (e.g., zobov; Neyrinck, 2008) to construct this continuous density field through Voronoi tessellation (or like Platen et al., 2007, use its dual, the Delaunay tessellation), where each galaxy or simulation particle is assigned all of the volume that is closer to it than to any other particle. The density for that whole region is then determined by dividing the particle’s mass by the volume, as if the matter had been spread evenly around the whole region.
@@ 3.1 Watershed segmentation
Figure 1: Density on a (200 $h^{-1}$ Mpc)${}^{2}$ region of a slice through the 4 $h^{-1}$ Mpc resolution simulation. Brighter green, towards yellow, indicates high density, and darker purple areas are of low density. The top panel shows the raw density field, the centre panel is annotated showing the segment boundary cells (black +), and the lower plot is annotated showing the void wall cells (black +) and over-dense excluded cells (yellow x).
@@ 3.2 Merging segments
Large voids are known to have a rich substructure (Tikhonov & Karachentsev, 2006; Sutter et al., 2014c); containing smaller voids and low-density structures. This can result in a slight density ridge inside a void, and the segmentation process will identify the areas on each side of such a ridge as separate watershed segments. However, it may be reasonable to consider both segments part of the same void. To this end, we perform a merging step after segmentation, which identifies segments that are separated by a wall with density below a chosen threshold and combines them into a single larger void. Any ridge with a density contrast below this threshold will be considered as part of the larger void surrounding it.
@@
where $\rho$ is the rest-mass density and $\bar{\rho}$ is its average over the spatial simulation slice. We choose $\delta=-0.8$ as our threshold, motivated by the density at which a top-hat model of an expanding under-density in an EdS universe experiences shell crossing (Sheth & van de Weygaert, 2004). Further, this same threshold and merging process is used in widely-used watershed void finders. It is a default for zobov (Neyrinck, 2008), a commonly used option in vide (Sutter et al., 2014a), and has been used in other implementations (e.g., Cautun et al., 2016).
@@ 3.5 Differences compared to particle-based methods
The walls that our void finder finds to separate voids must be a minimum of one grid cell thick. Physically, at our two resolutions, this means a thickness of 4 $h^{-1}$ Mpc and 12 $h^{-1}$ Mpc. . If finer detail could be resolved, it is likely that the wall is thinner than this, and therefore that the void should be larger than the volume we identify. This is particularly important for simulations with lower physical resolution. This is also expected to impact our void density profiles, as the matter is spread out over an entire grid cell. Since $N$-body simulations are capable of resolving smaller scales (particularly in walls, where the number density of particles is high), these will identify a more realistic wall structure.
@@ 4 Void Statistics
In this section, we outline the statistics of the sizes and shapes of voids in our simulations. We find a total of 30519 voids above the size cut-off in the (1024 $h^{-1}$ Mpc)${}^{3}$ volume of the 4 $h^{-1}$ Mpc resolution simulation, and 39206 in the (3072 $h^{-1}$ Mpc)${}^{3}$ volume at 12 $h^{-1}$ Mpc resolution.
@@ 4.1 Void size function
The void size function depends on the smoothing scale (Sheth & van de Weygaert, 2004), evident from the differences between the two panels in Fig. 2, so different simulations and observational datasets will also differ in their void size functions because of their various smoothing scales (see e.g., Nadathur & Hotchkiss, 2014; Schuster et al., 2023, for a study showing vastly different void size functions666 Specifically, in the isolated (no “segment merging”) catalogue of Schuster et al. (2023), comparable to our approach, VSFs for different resolutions differ, but their merged catalogue containing all subvoids and parent voids shows agreement for abundance of large voids at different resolutions. in different samples). Further, void size may be affected by the accelerated expansion resulting from a cosmological constant, so it is difficult to make a quantitative comparison between the void size function in this simulation without a cosmological constant and Newtonian $N$-body simulations that include a cosmological constant in the Friedmann equations that define the expansion of their Universe.

### v2

@@ Abstract
We apply and extend standard tools for void statistics to cosmological simulations that solve Einstein’s equations with numerical relativity (NR). We obtain a simulated void catalogue without Newtonian approximations, using a new watershed void finder which operates on fluid-based NR simulations produced with the Einstein Toolkit. We compare and contrast measures of void size and void fraction, and compare radial stacked density profiles to empirically-derived Hamaus–Sutter–Wandelt (HSW) density profiles and profiles based on distance to void boundaries. We recover statistics roughly consistent with Newtonian $N$-body simulations where such a comparison is meaningful. We study variation of dynamical spatial curvature and local expansion explicitly demonstrating the spatial fluctuations of these quantities in void regions. We find that voids in our simulations expand $\sim$10–30faster than the global average and the spatial curvature density parameter in the centre of voids reaches $\sim$60–80%.
@@
cosmology: theory – gravitation – large-scale structure of Universe – methods: statistical – methods: data analysis
@@ 1 Introduction
Observations present us with the immediate dichotomy that voids – the absence of matter – by their very nature can only be understood by directly observing the complex matter distribution within which they grow and interact. Observations need to be combined with modeling and numerical simulations to glean any understanding. The goal of this paper is to use numerical relativity to extend that modeling to a full general relativistic analysis of void statistics for the first time.
@@
Simulated void catalogues are typically produced by Newtonian $N$-body simulations, in which expansion is prescribed by scaling the simulation volume according to the Friedmann equation from an FLRW model. However, in GR matter and space-time curvature are non-linearly related. Consequently, regional expansion in the universe is determined by its local matter content, allowing over-dense and under-dense regions to have varying degrees of spatial curvature in any foliation. This may occur even if the average spatial curvature of the universe is zero (as in $\Lambda$CDM). This is particularly relevant for cosmic voids, which have a negative spatial curvature and expand at a greater rate than the over-dense sheets, filaments and knots of the cosmic web (Peebles, 1993, 2001; Macpherson et al., 2018). These concepts exist in standard cosmology. However, they are not explicitly implemented in the treatment of space-time in Newtonian simulations. In principle, variations of curvature and expansion in voids could be calculated from the Newtonian potential, but such an analysis has not yet been done.
A notable exception to traditional $N$-body methods is the heuristic AvERA scheme (Rácz et al., 2017) which replaces the Friedmann equation by an evolution code step that applies a volume-average to the scale factor governing the $N$-body simulation volume. This scheme has had phenomenological success in application to tensions in the integrated Sachs–Wolfe effect (Beck et al., 2018; Kovács et al., 2020, 2022), raising the question of whether such results can be replicated in numerical relativity.
Traditional $N$-body codes have been been significantly improved, most notably by gevolution (Adamek et al., 2016) which uses the weak-field limit of GR. In simplified setups, gevolution has been shown to agree with numerical relativity (Adamek et al., 2020, 2024). However, the weak-field scheme inherently precludes a full investigation of the potential for non-linear GR effects to significantly alter the chosen global background [^1] While a particular background cosmology must be chosen to set initial data in gevolution, in principle this can change during the evolution due to the absorption of homogeneous modes of the perturbations into the scale factor. However, this change in background is restricted to be small (see Section 5.3 of Adamek et al., 2016)..
Another alternative to evolving all degrees of freedom contained in the Einstein equations is to self-consistently remove degrees of freedom with likely small contributions in cosmology: (i) gravitational waves in the GRAMSES (Barrera-Hinojosa & Li, 2020a, b) code; (ii) vorticity and the magnetic Weyl curvature in quiet universes (Heinesen & Macpherson, 2022); and (iii) all of the above in silent universes (Bolejko, 2018a). Silent universe models have been used to seek solution to the Hubble tension (Bolejko, 2018b). However, the relationship between the Bolejko (2018b) solution and realistic initial conditions from the cosmic microwave background (CMB) epoch are not clear. In this work, we are interested in using numerical relativity to circumvent such limitations and study the fully non-linear problem.
@@ 1.2 Backreaction of inhomogeneities
Different types of backreaction have been classified in the well-known Buchert et al. (2020) (BMR) formalism, which provides a set of equations for the evolution of scalar averages of small scale structures. The first Buchert equation may be written as the energy density sum-rule:222 A cosmological constant backreaction density, $\Omega_{\Lambda}$, and additional backreaction terms are also found in general (BMR), but are not relevant for interpreting our simulations. The radiation density $\Omega_{\mathrm{R}}$ is not relevant except insofar as the ratio $\Omega_{\mathrm{R}}/\Omega_{\mathrm{M}}\simeq 0.3$ at last scattering, with significance for calibrations relative to CMB data. Our notation differs from that of BMR, who use $\Omega_{\mathcal{R}}$ in place of $\Omega_{\mathrm{K}}$, while not explicitly writing a radiation density. Positive values $\Omega_{\mathrm{K}}>0$ correspond to negative spatial curvature, as is the case for void domination.
$\Omega_{\mathrm{M}}+\Omega_{\mathrm{R}}+\Omega_{\mathrm{K}}+\Omega_{\mathcal{Q}}=1\,.$ (1)
Here $\Omega_{\mathrm{M}}$ and $\Omega_{\mathrm{R}}$ are analogous to matter and radiation energy densities in the Friedmann equation, but scales with respect to powers of an average volume scale $a_{\mathcal{D}}$, (cf. (11) below), not a background metric scale factor; $\Omega_{\mathrm{K}}$ is the average spatial curvature or kinetic curvature density; and $\Omega_{\mathcal{Q}}$ is the kinematical backreaction density.
Much attention has focused on the magnitude of $|\Omega_{\mathcal{Q}}|$, going back to the 2000s (Ishibashi & Wald, 2006). However, while a nonzero $|\Omega_{\mathcal{Q}}|$ is necessary for realistic cosmological backreaction,333 The imposition of periodic boundary condition in Newtonian cosmologies – the so-called torus condition – results in $\Omega_{\mathcal{Q}}$ necessarily vanishing in averages over the full global spatial hypersurfaces. The heuristic AvERA scheme (Rácz et al., 2017) explicitly breaks the assumptions of the Buchert–Ehlers (1997) theorem, and was criticised (Kaiser, 2017; Buchert, 2018) as a result. However, the quasilocal nature of gravitational mass in GR means that laws of mass conservation in GR may differ from their Newtonian counterparts. Thus the AvERA phenomenology may point to deeper fundamental questions. it may be small, as in the case of the timescape model (Wiltshire, 2007a, b, 2009; Duley et al., 2013). By contrast, a positive kinetic curvature term $\Omega_{\mathrm{K}}$ makes a very significant contribution to the late epoch energy budget (1) in several backreaction models, including those with an appreciable cosmological constant (Bolejko, 2018b). In the timescape model, $\Omega_{\mathrm{K}}$ has a direct physical interpretation as the quasilocal (i.e., regional) kinetic energy of expansion of voids – producing an apparent acceleration that does away with dark energy.
@@
Inhomogeneities are the source of local and quasilocal anisotropies in both GR and its weak field Newtonian limit. In fully relativistic cosmology regional anisotropies are not limited to a global FLRW background plus the spherical multipole expansion obtained by local Lorentz boosts in a $v/c$ power series. The differences have been classified via multipole moments as nonkinematic differential expansion (Bolejko et al., 2016; McKay & Wiltshire, 2016; Dam, 2016).
@@
The methods of Section 1.1 have enabled the first numerical investigations into the impact of backreaction (Bentivegna & Bruni, 2016; Macpherson et al., 2018, 2019) and cosmological observables (e.g. Giblin et al., 2017; Macpherson, 2023) in fully non-linear late-time cosmology. The characteristic features of quasilocal anisotropies expected in GR, as constrained by the initial matter power spectrum, are an important fundamental question. This paper quantifies the principal contribution to these small scale anisotropies – cosmic voids – for the first time.
@@ 2 Numerical Relativity Simulations
We evolve the simulations from the initial redshift through to $z\approx 0$ using the 3+1 Baumgarte–Shapiro–Shibata–Nakamura–Oohara–Kojima (BSSNOK or BSSN, Baumgarte & Shapiro, 1998; Shibata & Nakamura, 1995; Nakamura, Oohara & Kojima, 1987) formalism of NR. The final slice of the simulation is defined based on the predicted change in volume from the initial background EdS model. The space-time is evolved using the McLachlan (Brown et al., 2009) ML_BSSN thorn in the ET and the hydrodynamics is evolved using GRHydro (Baiotti et al., 2005). We approximate matter as dust, namely, as a pressureless [^6] GRHydro does not allow for an exactly zero pressure, so we choose $P\ll\rho$ using a polytropic equation of state (EOS), which is sufficient to approximate dust (Macpherson et al., 2017). Such an EOS has also been shown to be beneficial in analytic studies in inhomogeneous cosmology (Bolejko & Lasky, 2008). perfect fluid (under a continuous fluid approximation). We set the shift $\beta^{i}=0$ throughout and choose a harmonic-type slicing to evolve the lapse $\alpha$ according to (as in Macpherson et al., 2019) $\partial_{t}\alpha=-1/3\,\alpha^{2}K$ where $K$ is the trace of the extrinsic curvature. In the FLRW limit – or in the case of linear perturbations – this gauge coincides with a conformal-time slicing. When calculating the expansion and curvature for our void profiles, we calculate quantities intrinsic to the fluid flow in the simulation. These quantities (defined in Section 4.4.4) represent qualities of the frame that is everywhere orthogonal to the fluid 4-velocity, $u^{\mu}$. This fluid rest-frame is a more physically-motivated spatial slice than the particular slices we choose for the simulation itself. See Macpherson et al. (2017) and Macpherson et al. (2019) for further technical specifics of the simulation set up.
@@ 3 Watershed Void Finder
This method originates with the Watershed Void Finder (Platen, van de Weygaert & Jones, 2007) and zobov (Neyrinck, 2008). One popular watershed void finder based on zobov is vide (Sutter et al., 2015), which has been used in the study of both observational data (e.g., Sutter et al., 2012, 2014a) and simulations (e.g., Contarini et al., 2022).
Most void-finding software is designed to operate on sets of particle positions, either from observational data (positions of galaxies) or the output of $N$-body simulations (Platen et al., 2007; Neyrinck, 2008; Sutter et al., 2015). For watershed void-finding, it is necessary to translate these into a continuous density field upon which the watershed transform can be performed. It is typical (e.g., zobov; Neyrinck, 2008) to construct this continuous density field through Voronoi tessellation (or like Platen et al., 2007, use its dual, the Delaunay tessellation), where each galaxy or simulation particle is assigned all of the volume that is closer to it than to any other particle. The density for that whole region is then determined by dividing the particle’s mass by the volume, as if the matter had been spread evenly around the whole region.
@@ 3.1 Watershed segmentation
Figure 1: Density on a (200 $h^{-1}$ Mpc)2 region of a slice through the 4 $h^{-1}$ Mpc resolution simulation. Brighter green, towards yellow, indicates high density, and darker purple areas are of low density. The top panel shows the raw density field, the centre panel is annotated showing the segment boundary cells (black +), and the lower plot is annotated showing the void wall cells (black +) and over-dense excluded cells (yellow x).
@@ 3.2 Merging segments
Large voids are known to have a rich substructure (Tikhonov & Karachentsev, 2006; Sutter et al., 2014b); containing smaller voids and low-density structures. This can result in a slight density ridge inside a void, and the segmentation process will identify the areas on each side of such a ridge as separate watershed segments. However, it may be reasonable to consider both segments part of the same void. To this end, we perform a merging step after segmentation, which identifies segments that are separated by a wall with density below a chosen threshold and combines them into a single larger void. Any ridge with a density contrast below this threshold will be considered as part of the larger void surrounding it.
@@
where $\rho$ is the rest-mass density and $\bar{\rho}$ is its average over the spatial simulation slice. We choose $\delta=-0.8$ as our threshold, motivated by the density at which a top-hat model of an expanding under-density in an EdS universe experiences shell crossing (Sheth & van de Weygaert, 2004). Further, this same threshold and merging process is used in widely-used watershed void finders. It is a default for zobov (Neyrinck, 2008), a commonly used option in vide (Sutter et al., 2015), and has been used in other implementations (e.g., Cautun et al., 2016).
@@ 3.5 Differences compared to particle-based methods
The walls that our void finder finds to separate voids must be a minimum of one grid cell thick. Physically, at our two resolutions, this means a thickness of 4 $h^{-1}$ Mpc and 12 $h^{-1}$ Mpc. If finer detail could be resolved, it is likely that the wall is thinner than this, and therefore that the void should be larger than the volume we identify. This is particularly important for simulations with lower physical resolution. This is also expected to impact our void density profiles, as the matter is spread out over an entire grid cell. Since $N$-body simulations are capable of resolving smaller scales (particularly in walls, where the number density of particles is high), these will identify a more realistic wall structure.
@@ 4 Void Statistics
In this section, we outline the statistics of the sizes and shapes of voids in our simulations. We find a total of 30519 voids above the size cut-off in the (1024 $h^{-1}$ Mpc)3 volume of the 4 $h^{-1}$ Mpc resolution simulation, and 39206 in the (3072 $h^{-1}$ Mpc)3 volume at 12 $h^{-1}$ Mpc resolution.
@@ 4.1 Void size function
The void size function depends on the smoothing scale (Sheth & van de Weygaert, 2004), evident from the differences between the two panels in Fig. 2, so different simulations and observational datasets will also differ in their void size functions because of their various smoothing scales (see e.g., Nadathur & Hotchkiss, 2014; Schuster et al., 2023, for a study showing vastly different void size functions777 Specifically, in the isolated (no “segment merging”) catalogue of Schuster et al. (2023), comparable to our approach, VSFs for different resolutions differ, but their merged catalogue containing all subvoids and parent voids shows agreement for abundance of large voids at different resolutions. in different samples). Further, void size may be affected by the accelerated expansion resulting from a cosmological constant, so it is difficult to make a quantitative comparison between the void size function in this simulation without a cosmological constant and Newtonian $N$-body simulations that include a cosmological constant in the Friedmann equations that define the expansion of their Universe.